The EverySel package^{∗}

Martin Schröder

Friedrich-Humbert-Straße 124 D-28759 Bremen MS@Dream.HB.North.DE (Internet)

1996/05/30, L^{A}TEXed October 1, 1996

Abstract

This packages provides hooks into the NFSS-command \selectfont

     called \EverySelectfont and \AtNextSelectfont analogous to \AtBeginDocument.

Contents

1 Introduction 1

2 Usage 1

3 Options 2

4 Required packages 2

5 Acknowledgements 2

1 Introduction

This package provides the hooks \EverySelectfont and \AtNextSelectfont

whose arguments are executed just after L^{A}TEX has loaded a new font using \selectfont (which means that it will be executed after every font loaded via NFSS).

An example application would be a package for setting ragged text which needs to distiguinsh between monospaced and proportional fonts. Such a package exists: ragged2e.

2 Usage

\EverySelectfont \EverySelectfont{⟨code⟩} declares ⟨code⟩ that is saved internally and ex-

ecuted just after each \selectfont.

Warning: The ⟨code⟩ is saved globally; there is currently no way to remove it.

\AtNextSelectfont \AtNextSelectfont{⟨code⟩} declares ⟨code⟩ that is saved internally and

executed just after and only the next \selectfont.

Repeated use of the commands is permitted: the code in their argument is stored (and executed) in the order of their declarations.

The argument of \AtNextSelectfont is executed after the argument of

\EverySelectfont.

3 Options

The package has no options.

4 Required packages

The package requires no further packages.

5 Acknowledgements

David Carlisle provided the solution for my problems with \CheckCommand.

Index

Numbers written in italic refer to the page where the corresponding entry is described, the ones \underline{underlined} to the definition, the rest to the places where the entry is used.

AtNextSelectfont=  \subitem *+\AtNextSelectfont+, \usage{2}

                                    EverySelectfont=  \subitem *+\EverySelectfont+, \usage{1}

∗The version number of this file is v1.00, last revised 1996/05/30.

The name EverySel is a tribute to the 8 + 3 file-naming convention of certain “operating systems” and their “file systems”; strictly speaking it should be EverySelectfont.